# Equivariant Ensembles and Regularization for Reinforcement Learning in Map-based Path Planning

### v1

@@ I Introduction
To showcase the benefits of the equivariant ensembles and regularization, we evaluate their performance in a challenging, long-horizon, map-based planning application, the unmanned aerial vehicle (UAV) Coverage Path Planning (CPP) problem [7]. In this case study [^1] Video: https://youtu.be/L6NOdvU7n7s, the environment state can be represented as a map ([8]), and rotational symmetries can be exploited, as visualized in Figure 1. The results show that the ensemble makes the policy equivariant and that combining the ensemble and regularization improves performance significantly. We further show that regularization on the policy does not guarantee equivariance, which should be considered when regularizing the value estimate towards invariance.
@@
  • Implementation of ensembles and regularization in the long-horizon problem of UAV coverage path planning;
@@ II-B1 Fundamentals
in which $s^{\prime}$ and $a^{\prime}$ are the next state and action, respectively, which is the common notation adopted in the following. The Q-value determines the value of a specific action at the given state if following the policy afterward. Therefore, advantage is defined as
@@ VI-A Setup
To study the effect of equivariant ensembles and regularization, we trained agents on 10 maps with varying sizes from $32\times 32$ to $50\times 50$ shown in Figure 3(a). In each scenario encountered during training, one of the maps is randomly chosen, and a random target zone is generated. Since the range of possible target zones is vast, it is unlikely that an agent will ever encounter the same scenario between two episodes. Each agent is trained for 100M interaction steps with the environment with rollouts of 40K steps.
@@ VI-D Equivariance through Regularization
Baseline Augment Ensemble Regularize Ensemble + Regularize
[…]
$|\Delta\mathrm{V}(s)|$
[…]

### v2

@@ I Introduction
To showcase the benefits of the equivariant ensembles and regularization, we evaluate their performance in a challenging, long-horizon, map-based planning application, the unmanned aerial vehicle (UAV) Coverage Path Planning (CPP) problem [7]. In this case study, the environment state can be represented as a map ([8]), and rotational symmetries can be exploited, as visualized in Figure 1. The results show that the ensemble makes the policy equivariant and that combining the ensemble and regularization improves performance significantly. We further show that regularization on the policy does not guarantee equivariance, which should be considered when regularizing the value estimate towards invariance.
@@
  • Implementation [^1] Code: https://github.com/theilem/uavSim.git of ensembles and regularization in the long-horizon problem of UAV coverage path planning;
@@ II-B1 Fundamentals
in which $s^{\prime}$ and $a^{\prime}$ are the next state and action, respectively, which is the common notation adopted in the following. The Q-value determines the value of a specific action at the given state if following the policy afterward. Therefore, the advantage is defined as
@@ VI-A Setup
To study the effect of equivariant ensembles and regularization, we trained agents on 10 maps with varying sizes from $32\times 32$ to $50\times 50$ shown in Figure 4(a). In each scenario encountered during training, one of the maps is randomly chosen, and a random target zone is generated. Since the range of possible target zones is vast, it is unlikely that an agent will ever encounter the same scenario between two episodes. Each agent is trained using PPO [6] for 100M interaction steps with the environment with rollouts of 40K steps.
@@ VI-D Equivariance through Regularization
{tabu}
c—c——c—c—c—c—c    Baseline Augment Ensemble Regularize Ensemble + Regularize
[…]
@@
VII Discussion
In this paper, we proposed the usage of equivariant ensembles and regularization to exploit symmetries in MDPs. We proved that ensemble policies and value functions are respectively equivariant and invariant, providing theoretical soundness to value regularization. We demonstrated the benefits of the proposed approach in a map-based path planning case study, showing a decrease in sample complexity and an increase in performance and generalization.
This paper assumes that symmetries in the MDP are perfect. However, this assumption may not be valid in some real-world applications. In asymmetrical cases, our approach may still work as long as the asymmetry is observable to the agent; this we will study in future work. The additional computational overhead needs to be considered, especially when the group of transformations grows large, e.g., by adding the flip symmetries in the case study. In that case, we would need to investigate whether we can randomly sample a subset of the transformations, which could yield equivariance in expectation. Furthermore, in this specific case study, we saw the limits on out-of-distribution generalization. These limits could be overcome by training the agent on procedurally generated maps, for which the equivariant ensembles and regularization would play an important role.
In future work, we will also investigate the effect of equivariant ensembles and regularization for different problem settings, specifically for continuous action spaces. Further, to make it more broadly applicable, we will study the effects when applying different RL algorithms such as the off-policy SAC algorithm [18].